### v2

# Antidote: Post-fine-tuning Safety Alignment for Large Language Models against Harmful Fine-tuning

Abstract
Safety aligned Large Language Models (LLMs) are vulnerable to harmful fine-tuning attacks (Qi et al. 2023)– a few harmful data mixed in the fine-tuning dataset can break the LLMs’s safety alignment. Existing mitigation strategies include alignment stage solutions (Huang, Hu, and Liu 2024; Rosati et al. 2024a) and fine-tuning stage solutions (Huang et al. 2024; Mukhoti et al. 2023). However, our evaluation shows that both categories of defenses fail when some specific training hyper-parameters are chosen – a large learning rate or a large number of training epochs in the fine-tuning stage can easily invalidate the defense, which however, is necessary to guarantee finetune performance. To this end, we propose Antidote, a post-fine-tuning stage solution, which remains agnostic to the training hyper-parameters in the fine-tuning stage. Antidote relies on the philosophy that by removing the harmful parameters, the harmful model can be recovered from the harmful behaviors, regardless of how those harmful parameters are formed in the fine-tuning stage. With this philosophy, we introduce a one-shot pruning stage after harmful fine-tuning to remove the harmful weights that are responsible for the generation of harmful content. Despite its embarrassing simplicity, empirical results show that Antidote can reduce harmful score while maintaining accuracy on downstream tasks. Our project page is at https://huangtiansheng.github.io/Antidote_gh_page/
Introduction
Fine-tuning-as-a-service has become a new paradigm for Large Language Model (LLM) service with an increasing demand for personalized service delivery. Typically, fine-tuning data are uploaded by the users, and the service provider (e.g., OpenAI) finetunes a pre-trained LLM (foundation model), which is then served to meet the users’ customized need.
Before fine-tuning for user tasks, a pre-trained LLM is usually safety aligned to guarantee that the outputs of the LLM meet the safety preference, i.e., to refuse to generate harmful content even when the users trigger them to do so. However, recent studies (Qi et al. 2023; Yang et al. 2023; Zhan et al. 2023) show that a few harmful data mixed in the fine-tuning dataset can trigger the model to forget the alignment knowledge it learned previously –it no longer uses refusal response when users submit a harmful prompt.
Figure 1: Antidote with a three-stage pipeline, i.e., i) safety alignment, ii) user fine-tuning, iii) one-shot pruning. While existing works mostly focus on optimization in the first stage, e.g., (Huang, Hu, and Liu 2024; Rosati et al. 2024a) or the second stage (Huang et al. 2024; Mukhoti et al. 2023), Antidote utilizes the post-fine-tuning stage to prune the harmful weights to recover the model from harmful behaviors.
Existing mitigation strategies can be broadly categorized into two categories, i.e., alignment stage defense and user fine-tuning stage defense. The first category is concerned with how to improve the large language model’s immunization towards the harmful fine-tuning data in the alignment stage. For example, (Huang, Hu, and Liu 2024) add artificial perturbation in the alignment stage to simulate the harmful embedding drift in the fine-tuning stage, and utilizes a mini-max optimization to enforce the model to be immune to the perturbation. (Rosati et al. 2024b) utilize a representation noising technique to degrade the representation distribution of the harmful data to a random Gaussian noise, such that the harmful content generation is more difficult to learn by harmful fine-tuning data. For fine-tuning-stage mitigation, the core idea is to mitigate the forgetting of the alignment knowledge but also to learn the knowledge for the users’ tasks. To achieve this goal, (Huang et al. 2024) separate the fine-tuning stage into two states, which alternatively optimize the alignment and fine-tuning dataset. (Mukhoti et al. 2023) add a regularization to constrain drift in feature space to mitigate the forgetting of alignment knowledge. However, our empirical evaluation in Section III reveals a common weakness of the existing defense methods – a small learning rate and a small number of epochs in the fine-tuning stage are required to guarantee their effectiveness. This requirement can be detrimental to downstream tasks’s performance because some fine-tuning tasks require a larger learning rate and longer training epochs to guarantee learning performance.
To this end, we in this paper aim to answer the following research question:
Is there a safety alignment that can be insensitive to the hyper-parameters for the fine-tuning stage?
Guiding by this urge, we propose Antidote, a safety alignment solution that realigns the model after the fine-tuning stage has been fully completed. The design of Antidote is agnostic to how the fine-tuning is done – it relies on the philosophy that by removing the harmful parameters, the harmful model can be recovered from the harmful behaviors, regardless of how those harmful parameters are formed in the fine-tuning stage. Empirically, we show that Antidote respectively reduces the harmful score by up-to 17.8% (compared to SFT without defense) while maintaining the same level of fine-tuning accuracy (by up-to 1.83% accuracy loss).
To the end, we summarize our contribution as follows:
  • We evaluate the existing solutions for harmful fine-tuning. We show that existing solutions are highly sensitive to the training hyper-parameters in the fine-tuning stage, which we name hyper-parameter sensitive issue.
  • To fix this issue, we propose Antidote, a post-fine-tuning realignment solution that remains agnostic towards the training details in the fine-tuning stage.
  • Comprehensive experiments on four downstream tasks and different attack settings are conducted to verify the effectiveness of the proposed method.
Related work
Safety alignment. Safety alignment is about how to align a large language model such that its outputs are aligned with humans’ values. Representative techniques are RLHF (Ouyang et al. 2022) and its variants (Dai et al. 2023; Bai et al. 2022; Wu et al. 2023; Dong et al. 2023; Rafailov et al. 2023; Yuan et al. 2023). Most recently, there are alternative solutions focusing on augmenting the alignment data, e.g., (Liu, Sferrazza, and Abbeel 2023; Liu et al. 2023; Ye et al. 2023).
Harmful fine-tuning. (Qi et al. 2023; Yang et al. 2023; Zhan et al. 2023) show that LLMs aligned by RLHF or SFT (supervised fine-tuning) can be jail-broken after fine-tuning on explicit/implicit harmful data, and several mechanism studies (Leong et al. 2024; Wei et al. 2024; Peng et al. 2024; Jain et al. 2024) are conducted to analyze the problem. Existing solutions for harmful fine-tuning issues can be categorized into two main categories. The first category is alignment stage solutions, which study how to improve the model’s immunization ability to the fine-tuning by modifying training procedure in alignment stage. Exemplar literature are Vaccine (Huang, Hu, and Liu 2024) and RepNoise (Rosati et al. 2024b, a). Specifically, Vaccine vaccinates the model by adding embedding perturbation in the alignment stage, and RepNoise improve the robustness by enforcing the representation of the harmful data to be a random Gaussian noise.A concurrent research (Tamirisa et al. 2024) utilize meta learning to immunize the model. The second category is fine-tuning-stage solutions (Mukhoti et al. 2023; Zong et al. 2024; Huang et al. 2024; Wang et al. 2024; Lyu et al. 2024)., which study how to avoid forgetting the alignment knowledge while also learning the fine-tuning knowledge by modifying training procedure in user fine-tuning stage. Specifically, LDIFS (Mukhoti et al. 2023) introduces a regularizer to enforce the iterate’s embedding to be in close proximity to that of the aligned model. VLGuard(Zong et al. 2024) mixes the alignment data with the user data to train in the user fine-tuning stage. Lisa(Huang et al. 2024) alternatively optimizes over the alignment data and the user fine-tuning data and use proximal regularizer to enforce proximity between iterates. Recently, there are also more advanced attack (He, Xia, and Henderson 2024; Halawi et al. 2024) in this setting.
Model sparsification. Since (Frankle and Carbin 2018), model sparsification for deep learning models has been extensively studied. The core research problem for model sparsity literature is to score the weights coordinates according to their importance. The coordinates with small importance score are then removed to compress the model. Specifically for large language model, (Frantar and Alistarh 2023) propose SparseGPT, which forms the importance score by solving a layer-wise reconstruction problem. (Sun et al. 2023) propose Wanda score, which utilize joint weights/activation metrics to measure the coordinate importance. On top of Wanda, (Yin et al. 2023) propose layer-wise sparsity, which further improve the model compression ratio. We in this paper borrow importance score from the model sparsification literature to identify and remove harmful parameters.
We acknowledge that there are a few concurrent post-fine-tuning stage defenses, which also aim at purifying the model after fine-tuning completes. For example, (Hsu et al. 2024) project the harmful gradient updates to a aligned subspace. (Yi et al. 2024) realign model via subspace-oriented model fusion. (Tong et al. 2024) realign by self-contrastive decoding. It is possible that these existing solutions can also be insensitive the training hyper-parameters in fine-tuning stage due to the good nature of post-fine-tuning solutions. However, prior to this study, there is no systematical study on the hyper-parameter sensitivity issue, which highlights the significance of post-fine-tuning stage defense.
Preliminaries
Threat Model and Assumptions
A typical fine-tuning-as-a-service pipeline. is considered in (Qi et al. 2023), where the users upload a few fine-tuning data to the service provider. On behalf of users, the service provider fine-tunes the aligned pre-trained model on this dataset, and the finetuned model is deployed by the service provider to deliver personalized service. Because the model is deployed in the service provider’s server and the answer to user prompts is delivered by the service provider’s API, the service provider has an obligation to ensure the answer is harmless. Otherwise, the service provider might face governance issues (Reuel et al. 2024) or lawsuit  [^1] Laws have been enforced, e.g., SB-1047 in California..
Following (Rosati et al. 2024a; Tamirisa et al. 2024), we assume the service provider hosts a harmful dataset $\mathcal{D}_{realign}$ (containing harmful prompt-harmful answer pairs), which we use to perform post-fine-tuning stage re-alignment. We henceforth refer to this dataset as re-alignment dataset for clearness. Following (Rosati et al. 2024a; Huang, Hu, and Liu 2024; Hsu et al. 2024; Zong et al. 2024), we assume the service provider maintains a safety alignment dataset $\mathcal{D}_{align}$ (containing harmful prompt-safe answer pairs).
Revisiting Existing Defenses
In this subsection, we evaluate the existing safety alignment for harmful fine-tuning issue. We choose two representative alignment-stage solutions (Huang, Hu, and Liu 2024; Rosati et al. 2024a) and two fine-tuning-stage solutions (Huang et al. 2024; Mukhoti et al. 2023) as demonstration.
Figure 2: Harmful score and finetune accuracy with different learning rates in the fine-tuning stage. Here we fix fine-tuning epochs to 20. Harmful score and finetune accuracy with different learning rates in the fine-tuning stage. Here we fix fine-tuning epochs to 20.
Existing defenses fail with a large learning rate in fine-tuning stage. We adjust the learning rate in the user fine-tuning stage and show how existing methods perform in Figure 2. Our results show that both alignment-stage defenses (Vaccine and RepNoise) and fine-tuning-stage defenses (Lisa and LDIFS) tend to have larger harmful scores when learning rate is large. We now explain the reason for their failures. i) For alignment stage solutions, the core idea of defense is to strengthen the aligned model’s robustness towards harmful data in the later fine-tuning stage. The reason for a degraded performance is that a larger learning rate in fine-tuning stage can make it easier to subvert the model robustness induced before fine-tuning (the same phenomenon and explanation are given in (Rosati et al. 2024a)). ii) For fine-tuning stage defenses, the core idea is to introduce a regularizer in the fine-tuning stage to enforce the fine-tuning iterate in proximity to the aligned model. These solutions also suffer degraded performance because a larger learning rate may drift the iterates far away from the aligned model, resulting in the model failing to converge near the aligned model.
Figure 3: Harmful score and finetune accuracy with different fine-tuning epochs in the user fine-tuning stage. Here we fix fine-tuning learning rate to 1e-5.
Existing defenses fail with a large number of fine-tuning epochs. We adjust the number of fine-tuning epochs in the user fine-tuning stage and show the results in Table 3. Similar to the learning rate, a larger number of fine-tuning epochs tends to enlarge the harmful score and break the defense. The reasons for their failure are similar to that induced by the large learning rate, i.e., i) strengthened alignment can still be jail-broken with more training epochs, and ii) More fine-tuning epochs induce more drift towards the aligned iterate.
A sufficiently large learning rate and finetune epochs are necessary. However, as shown in Table 2 and 4, a sufficiently large learning rate is necessary to guarantee good fine-tune accuracy, which indicates that state-of-the-art solutions fall short and need renovation.
We refer to the common weakness of these defenses as hyper-parameter sensitivity issue, which restricts the general usage of the alignment solutions.
Methodology
In order to counter the hyper-parameters sensitivity issue in fine-tuning stage, we propose a post-fine-tuning safety alignment that remains agnostic to the exact training setting in fine-tuning. The high-level idea is to remove the harmful parameters in the model after the model has corrupted with fine-tuning. The method is agnostic to the hyper-parameter in fine-tuning stage because in principle the harmful parameters can anyway be deactivated regardless of how they form in the fine-tuning stage. We refer to Figure 4 for a system overview.
Figure 4: Detailed procedure of Antidote. On Stage III after model has been fine-tuned, Antidote extracts the importance masks over realignment dataset. Then this mask is applied to purify the harmful fine-tuned model.
Algorithm 1 Antidote: a post-fine-tuning safety alignment
0:  Mask ratio, $\alpha$;Alignment dataset, $D_{align}$; Fine-tuning dataset, $D_{ft}$; Re-alignment dataset, $\mathcal{D}_{realign}$;
0:  The re-aligned model $\tilde{\bm{w}}$ ready for deployment.
  — — STAGE I (alignment) begins here — —
  $\check{\bm{w}}$ = SafetyAlignment($\bm{w}$) on $D_{align}$
  — — STAGE II (fine-tuning) begins here — —
  $\hat{\bm{w}}$ = Hamrfulfine-tuning($\check{\bm{w}}$) on $D_{ft}$
  — STAGE III (one-shot pruning) begins here
  Calculate importance score $h(\hat{\bm{w}},\mathcal{D}_{\text{realign}})$ with Eq. (1)
  $\hat{\bm{m}}$ = ArgTopKα($h(\hat{\bm{w}},\mathcal{D}_{realign})$)
  $\tilde{\bm{w}}=(\bm{1}-\tilde{\bm{m}})\odot\hat{\bm{w}}$
To achieve the defense goal, we first need to identify the important parameters over a particular dataset using Wanda score. The Wanda score (Sun et al. 2023) measures the importance score of parameters given a dataset $\mathcal{D}$, as follows:
$[h(\bm{w},\mathcal{D})]_{j}=\frac{1}{|\mathcal{D}|}\sum_{\bm{X}\in\mathcal{D}}|\bm{w}_{j}|\cdot\|\bm{X}_{j}\|_{2}$ (1)
where $[\cdot]_{j}$ retrieve the j-th element of the vector, $\bm{X}_{j}$ is the input (or activation) associated with the $j$-th weight coordinate, and $\bm{X}$ represents a data point in a given dataset $\mathcal{D}$.
To recognize the harmful parameters, one intuitive idea is to extract the topk most important mask (harmful mask) on the re-alignment dataset, indicating the most important parameters for harmful content generation, as follows.
$\hat{\bm{m}}=\text{ArgTopK}_{\alpha}(h(\hat{\bm{w}},\mathcal{D}_{realign}))$ (2)
Given the harmful mask $\tilde{\bm{m}}$ and the weights after fine-tuning $\hat{\bm{w}}$, The pruning operation is as follows:
$\tilde{\bm{w}}=(\bm{1}-\tilde{\bm{m}})\odot\hat{\bm{w}}$ (3)
where $\tilde{\bm{w}}$ is the re-aligned model that is ready for deployment.
In summary, at stage I, we align the model on the alignment dataset using a vanilla SFT solution. At stage II, we finetune the model on the user fine-tuning dataset with SFT, and obtain the harmful mask, which captures the important parameters for harmful data. Finally, at the post-fine-tuning stage, we apply the pruning mask to the model and recover it from the harmful behavior. See Algorithm 1 for full procedure.
Table 1: Harmful score and finetune accuracy under different harmful ratio. Other settings are default.
Methods Harmful score Finetune accuacy
clean p=0.05 p=0.1 p=0.2 p=0.5 Average clean p=0.05 p=0.1 p=0.2 p=0.5 Average
SFT 52.30 76.70 79.00 79.40 80.20 73.52 95.87 95.18 95.07 95.18 93.69 95.00
Repnoise 42.40 79.20 79.50 77.90 82.60 72.32 95.07 94.84 94.84 94.38 94.61 94.75
Vaccine 44.80 80.20 80.00 81.50 81.90 73.68 95.53 95.53 94.04 95.18 94.04 94.86
Lisa 53.00 60.90 64.80 68.20 72.10 63.80 93.92 93.69 93.58 93.23 91.17 93.12
LDIFS 51.70 67.70 68.80 72.30 71.80 66.46 93.46 93.23 93.69 93.23 94.04 93.53
Antidote 52.90 61.20 61.20 64.60 64.50 60.88 93.58 93.46 93.12 93.35 91.74 93.05
Experiments
Setup
Model and Datasets. We use three mainstream pre-trained models, i.e., Llama2-7B, Mistral-7B and Gemma-7B for evaluations. In the default setting, we use Llama2-7B as the backbone. We consider three datasets associated with harmful data. The first dataset is an alignment dataset, which contains alignment data (i.e., data paired with harmful prompt-safe answers). The second is fine-tuning the dataset. This dataset is mixed with $p$ (percentage) of harmful data (paired with harmful prompt-harmful answer) and $1-p$(percentage) of downstream data (e.g., SST2, GSM8K, etc). The last one is a re-alignment dataset, which is solely constituted by harmful data. The alignment data are sampled from BeaverTails (Ji et al. 2023) with the label is_safe=True. The harmful data in fine-tuning dataset and realignment are also sampled from BeaverTails (Ji et al. 2023) with is_safe=False, but the harmful data in those two datasets are different. For fine-tuning tasks, we consider four different datasets, i.e., SST2, AGNEWS, GSM8K and AlpacaEval. We discuss how to integrate and evaluate these tasks in supplementary materials.
Metrics. Following (Huang et al. 2024), we use two metrics for evaluation.
  • Finetune Accuracy (FA). It is Top-1 accuracy of the model over the fine-tuning task’s test dataset.
  • Harmful Score (HS). We use the moderation model from (Ji et al. 2023) to flag the model output given unseen malicious instructions. Harmful score is the ratio of the flagged unsafe output.
To calculate the harmful score, we sample 1000 harmful instructions from BeaverTails (Ji et al. 2023). To calculate finetune accuracy, we sample 872, 1000, 1000, and 122 samples from the corresponding fine-tuning testing dataset.
Baselines. We mainly consider five baselines in evaluation. SFT is utilized to supervised fine-tuning on both the alignment and fine-tuning stages. Two representative alignment-stage solutions, i.e., Vaccine (Huang, Hu, and Liu 2024) and RepNoise (Rosati et al. 2024a) modify the alignment stage while keeping the fine-tuning stage optimization as SFT. Two representative fine-tuning-stage solutions, i.e., Lisa(Huang et al. 2024) and LDIFS (Mukhoti et al. 2023) modify the fine-tuning stage while keeping the alignment stage optimization as SFT.
Training Details and Hyper-parameters. We follow (Huang, Hu, and Liu 2024) to utilize LoRA (Hu et al. 2021) for alignment and fine-tuning. The rank of the adaptor is 256 for both tasks. For alignment, we use 5000 safety samples and an AdamW optimizer with a learning rate of 1e-3. We train the alignment data for 20 epochs. For fine-tuning, we use $n$ samples, among which $p$ (percentage) of samples are harmful data, an AdamW optimizer with a learning rate of $lr$ is used, and we train for $ep$ epochs. The default setting is $n=5000$, $p=0.2$, $lr=1e-4$ and $ep=20$, and the default dataset is SST2 unless otherwise specified. For hyper-parameters for Antidote, in default, we set the mask ratio to be $\alpha=0.2$ (specially, $\alpha=0.05$ for GSM8K) and we sample 2000 harmful samples to form the re-alignment dataset used by Antidote. See for a setting for hyper-parameters for baselines. All experiments are done with an H100.
Table 2: Performance under different number of fine-tuning samples. While Lisa achieves the smallest average harmful score, its finetune accuracy is unacceptably low.
Methods Harmful score Finetune accuacy
n=100 n=1000 n=2000 n=3000 n=5000 Average n=100 n=1000 n=2000 n=3000 n=5000 Average
SFT 65.50 76.90 77.80 80.70 79.40 76.06 92.20 94.72 94.27 94.50 95.18 94.17
Repnoise 66.50 77.60 78.80 78.60 77.90 75.88 89.45 92.66 93.69 94.72 94.38 92.98
Vaccine 66.40 79.00 78.60 81.10 81.50 77.32 90.48 93.92 94.95 95.30 95.18 93.97
Lisa 52.80 52.40 54.00 64.30 68.20 58.34 26.72 33.72 49.54 91.17 93.23 58.88
LDIFS 55.70 64.60 67.10 68.90 72.30 65.72 87.73 91.17 92.32 92.43 93.23 91.38
Antidote 57.00 60.70 62.80 61.70 64.60 61.36 90.02 92.43 93.12 93.00 93.35 92.38
Main Results
Robustness to harmful ratio. We show in Table 1 how different methods perform when different ratios of harmful data are mixed into the fine-tuning data. Our results indicate that Antidote is able to achieve the lowest harmful score in most settings harmful ratio – it achieves a remarkable 11.56% reduction of average harmful score compared to SFT with a marginal 1.45% loss of average finetune accuracy. It is also notable that Antidote is able to achieve consistently good defense performance with no clear trend of performance degradation when the harmful ratio is high. In contrast, all the other methods tend to lose effectiveness when the harmful ratio is high (for example, Lisa has an 11.2% increase of harmful score from $p=0.05$ to $p=0.5$). The advantage of Antidote comes from the post-fine-tuning design, which remains agnostic of how different ratios of harmful data originally aligned model, and is not sensitive to how the model is going to drift from the aligned model produced by the previous alignment stage. Of note, the two alignment stage solutions do not seem to work well in all the settings. We will delay the analysis of their failure in the later learning rate experiment.
Table 3: Harmful score and finetune accuracy under different learning rate. The dataset is GSM8K and other settings are default.
Methods Harmful score Finetune accuacy
lr=1e-7 lr=1e-6 lr=1e-5 lr=1e-4 lr=1e-3 Average lr=1e-7 lr=1e-6 lr=1e-5 lr=1e-4 lr=1e-3 Average
SFT 52.80 70.30 80.10 77.80 79.80 72.16 4.30 14.00 23.10 21.90 23.30 17.32
Repnoise 52.50 70.10 79.00 80.20 75.50 71.46 4.80 12.60 24.90 23.50 24.70 18.10
Vaccine 46.50 66.00 79.40 80.60 77.50 70.00 1.80 10.90 25.50 24.20 25.80 17.64
Lisa 52.30 55.00 64.40 73.20 77.30 64.44 4.00 5.70 13.60 21.90 24.70 13.98
LDIFS 53.20 56.10 59.00 68.50 78.50 63.06 4.00 4.80 5.40 6.10 14.10 6.88
Antidote 53.50 61.80 65.60 65.30 68.80 63.00 4.10 11.20 17.50 16.10 20.40 13.86
Table 4: Evaluation under different training epochs. The dataset is GSM8K and other settings are default.
Methods Harmful score Finetune accuacy
ep=1 ep=5 ep=10 ep=20 ep=40 Average ep=1 ep=5 ep=10 ep=20 ep=40 Average
SFT 76.50 78.90 79.90 77.80 78.70 78.36 21.00 25.80 26.50 21.90 24.60 23.96
Repnoise 76.30 79.50 79.00 80.20 80.80 79.16 19.70 26.20 26.10 23.50 22.70 23.64
Vaccine 75.80 82.10 79.60 80.60 80.40 79.70 20.40 26.00 25.10 24.20 22.60 23.66
Lisa 55.40 54.80 71.50 73.20 75.00 65.98 4.50 4.50 21.70 21.90 24.40 15.40
LDIFS 56.70 61.50 64.90 68.50 72.40 64.80 4.90 5.00 5.70 6.10 6.10 5.56
Antidote 61.50 66.80 66.60 65.30 63.60 64.76 13.60 17.80 19.80 16.10 13.90 16.24
Robustness to fine-tuning samples. We show in Table 2 how different fine-tuning samples used in the fine-tuning stage will affect the practical defense performance. Our results indicate Antidote again achieves the best defense performance among the baselines with a remarkable 13.42% reduction of the harmful score. Antidote is again the only defense that is universally robust to different sample number settings.
Robustness to learning rate in fine-tuning. In Table 3, we adjust the learning rate in the fine-tuning stage to see its impact on defense performance. We use GSM8K for this evaluation to provide more diversified statistical data. Our results show that Antidote reduces 6.56% average harmful score with a marginal 0.38% average finetune accuracy drop. Although the harmful score reduction is less competitive compared to two fine-tuning-stage defenses Lisa and LDIFS (which respectively achieve 7.72% and 9.50% harmful score reduction), we note that their performance gain comes with a drastic reduction of finetune accuracy. Moreover, the result when $lr=1e-3$ also coincides with the finding in our motivation section, that both Lisa and LDIFS suffer from hyper-parameter sensitivity. In contrast, Antidote is less susceptible to the exact setting of learning rate.
Robustness to training epochs in fine-tuning. In Table 4, we adjust the number of training epochs in the fine-tuning stage to see its impact on the defense performance. GSM8K is used for this evaluation to provide more diversified statistical data. The results show that while other defenses tend to have a larger harmful score when fine-tuning epochs are larger, this is not obvious for Antidote (specifically, harmful score is 6.3% lower for Antidote from $ep=10$ to $ep=40$). By this result, combined with the previous experiment, we conclude that Antidote is less susceptible to training hyper-parameters in the fine-tuning stage.
Generalizations to datasets. In Table 5, we show the evaluation results for different datasets. Our results confirm that Antidote can be generalized to different fine-tuning tasks. On average, we show that Antidote reduce the harmful score by 11.75% with 3.08% of finetune accuracy loss. Here we did not specifically tune the mask ratio $\alpha$ for each dataset, and we will discuss later that the tradeoff between harmful score and finetune accuracy can be adjusted by this key hyper-parameter.
Table 5: Evaluation on different datasets.
Methods SST2 AGNEWS GSM8K AlpacaEval Average
HS FA HS FA HS FA HS FA HS FA
SFT 79.40 95.18 79.60 92.70 77.80 21.90 73.80 43.27 77.65 63.26
Repnoise 77.90 94.38 82.30 92.20 80.20 23.50 73.50 42.00 78.90 63.14
Vaccine 81.50 95.18 81.10 93.00 80.60 24.20 73.40 40.10 79.15 63.12
Lisa 68.20 93.23 74.80 90.80 73.20 21.90 65.20 39.90 72.45 61.92
LDIFS 72.30 93.23 69.60 87.10 68.50 6.10 66.60 39.81 69.25 56.56
Antidote 64.60 93.35 69.50 88.00 65.30 16.10 60.50 41.83 64.98 59.82
Generalizations to models. We show in Table 6 how different methods perform on different LLMs. Our results indicate that Antidote can be generalized to different model architectures. For Llama2-7B, Mistral-7B, and Gemma-7B, Antidote respectively achieve 11.6%, 20.0%, 22.5% reduction of harmful score with a minor reduction of 1.49%, 0.92%, and 1.72% finetune accuracy. Particularly, in terms of finetune accuracy, our results coincide with the bench-marking results that the rank of language ability of three models is Gemma-7B$>$Mistral-7B$>$Llama2-7B. In terms of reducing harmful score, our results seem to indicate that Antidote is more effective when the backbone model is stronger.
Table 6: Evaluation on different models.
Methods Llama2-7B Mistral-7B Gemma-7b Average
HS FA HS FA HS FA HS FA
SFT 79.40 95.18 80.30 95.99 80.90 96.22 80.20 95.80
Repnoise 77.90 94.38 79.00 94.95 80.70 88.76 79.20 92.70
Vaccine 81.50 95.18 80.60 94.04 79.10 94.72 80.40 94.65
Lisa 68.20 93.23 65.30 95.07 75.40 96.22 69.63 94.84
LDIFS 72.30 93.23 69.50 92.09 72.70 93.35 71.50 92.89
Antidote 64.60 93.35 64.80 94.95 59.40 94.04 62.93 94.11
Statistical and System Evaluation
Figure 5: Harmful embedding drift (HED) under different learning rate and epochs in fine-tuning stage. Antidote obtains a relatively small HED.
Harmful embedding drift. To justify the reason why Antidote is able to recover the model even with a large learning rate and training epoch, we follow (Huang, Hu, and Liu 2024) to measure the harmful embedding drift (HED), which tracks the L2 norm of the difference between the hidden embedding of the aligned model and that of the finetuned model over the same alignment data. We show in Figure 5 how different learning rates and training epochs affect this statistic. As shown, Antidote maintains the HED on a small scale, while the HED of other mitigation strategies escalates with the growth of learning rate and training epochs. Particularly, note that Antidote and SFT share the same two identical processes (and therefore the same HED) in their first two stages, but after the one-shot pruning in the post-fine-tuning stage, the HED of Antidote is significantly lower. This justifies that pruning out the identified harmful parameters can indeed recover the alignment knowledge preserved in the model.
In summary, our finding justifies that Antidote can minimize the hidden embedding drift over the alignment data by pruning some specific harmful coordinates, and thereby mitigating the harmful embedding drift issue mentioned in (Huang, Hu, and Liu 2024).
System performance. We then measure the system performance (i.e., clock time, GPU memory usage) of different solutions in Table 7. Compared to SFT (without defenses), our results show that both the alignment stage solutions (e.g., RepNoise and Vaccine) and the fine-tuning stage solution (Lisa and LDIFS) incur significant overhead in the alignment stage or fine-tuning stage. For example, Vaccine and RepNoise require over 2x clock time for alignment. and over 1.7x GPU memory consumption. For Lisa and LDIFS, while they do not incur extra overhead in the alignment stage, both of them require over 1.6x GPU memory, and LDIFS requires 1.5x clock time in the fine-tuning stage. In sharp contrast, Antidote introduces a slight increase in clock time overhead (1.02x clock time) and the same GPU memory usage for the whole pipeline. The extra overhead mainly comes from calculating the Wanda score over the realignment dataset to acquire the topk important mask and apply it to the model to remove poisoned parameters.
Table 7: System evaluation for different methods. Antidote introduces extra overhead in post-fine-tuning stage mainly due to Wanda score calculation.
Methods Clock time (hour) GPU Memory (GB)
Alignment Fine-tuning Post-FT Sum Alignment Fine-tuning Post-FT Max
SFT 0.92 (1x) 0.78 (1x) 0 1.70 (1x) 35.45 (1x) 33.06 (1x) 0 35.45 (1x)
Repnoise 1.97 (2.14x) 0.78 (1x) 0 2.75 (1.62x) 75.26 (2.12x) 33.06 (1x) 0 75.26 (2.12x)
Vaccine 1.84 (2x) 0.78 (1x) 0 2.63 (1.54x) 56.46 (1.71x) 33.06 (1x) 0 56.46 (1.71x)
Lisa 0.92 (2.14x) 0.80 (1.03x) 0 1.72 (1.01x) 35.45 (1x) 52.95 (1.60x) 0 52.95 (1.49x)
LDIFS 0.92 (2.14x) 1.19 (1.53x) 0 2.11 (1.24x) 35.45 (1x) 64.53 (1.95x) 0 64.53 (1.82x)
Antidote 0.92 (1x) 0.78 (1x) 0.04 1.78 (1.02x) 35.45 (1x) 33.06 (1x) 22.35 35.45 (1x)
Hyper-parameters Analysis and Ablation Study
Impact of mask ratio $\alpha$. We next show how Antidote performs given different mask ratios as its hyper-parameter in Table 8. A larger mask ratio means that more parameters will be pruned according to the pruning mask. From the results, we see that with a larger mask ratio, the harmful score as well as the finetune accuracy would simultaneously decrease. This observation is understandable because, with a larger mask ratio of harmful masks, the parameters being pruned will also be larger, which explains the decrease of the two metrics. Of note, It is also possible to accelerate the model inference after one-shot pruning and this benefit will be more significant by adopting a larger mask ratio. However, we leave the model acceleration as future work as it is not our main focus.
Table 8: Evaluation of Antidote under different mask ratio $\alpha$.
$\alpha$=0.01 $\alpha$=0.05 $\alpha$=0.1 $\alpha$=0.15 $\alpha$=0.2 $\alpha$=0.25
HS 73.60 68.70 64.60 58.90 58.40 57.00
FA 94.95 94.50 93.35 91.06 86.58 80.05
Necessity of re-alignment dataset. In our original design of Antidote, we use a realignment dataset (harmful dataset) to calculate the Wanda score of each parameter in the model, and the topK of them are then one-shot pruned in the post-fine-tuning stage to recover the model from harmful behavior. In Table 9, we show how the defense performance will become by replacing the realignment dataset with the fine-tuning dataset or benign dataset (fine-tuning dataset after precluding harmful data). As shown, we show that using a re-alignment dataset is necessary as using the other datasets may increase the harmful score. Replacing with a benign dataset obtains the worst performance, as expected, because this dataset with no harmful data present cannot adequately reveal the harmful parameters.
Table 9: Evaluation of Antidote with different ways to calculate Wanda score under different poison ratio $p$. Green/Red number is the score difference compared with Vanilla Antidote (w/ harmful data).
p=0.05 p=0.1 p=0.2 p=0.5 Average
HS (w/ harmful data) 63.10 68.30 68.80 69.20 67.35
HS (w/ fine-tuning data) 63.30 (+0.20) 69.80 (+1.50) 68.50 (-0.30) 70.50 (+1.30) 68.03 (+0.68)
HS (w/ benign data) 63.80 (+0.70) 69.70 (+1.40) 69.20 (+0.40) 71.20 (+2.00) 68.48 (+1.13)
Impact of the size of realignment dataset. Per Eq. (1), we calculate the harmful Wanda score with the average statistics over the realignment dataset (harmful dataset). It is interesting to see how different numbers of harmful samples included in the re-alignment dataset can affect defense performance. Our results in Table 10 show that the general trend is that a larger number of harmful samples are preferable in terms of reducing harmful score. This is understandable because a larger number of harmful samples can better reflect the real harmful distribution therefore resulting in a more precise identification of harmful parameters. Another finding is that 1k samples of harmful data seem to be performing well, and the benefit of further increasing the sample number diminishes. As collecting 1k harmful samples is not too restrictive [^2] We have much larger scale harmful datasets available today. For example, BeaverTail (Ji et al. 2023) contains 330k harmful samples., this experiment validates the feasibility of Antidote.
Table 10: Harmful score (HS) of Antidote with different number of harmful samples in realignment dataset. When $|\mathcal{D}_{realign}|=0$, the wanda score of each coordinate reduces to its weight magnitude.
$|\mathcal{D}_{realign}|$ 0 5 10 100 1k 2k 5k
HS 72.1 70.60 70.30 70.20 69.30 69.20 69.40
(0) (-1.5) (-1.8) (-1.9) (-2.8) (-2.9) (-2.7)
Extensions
As Antidote is a post-fine-tuning stage solution, it is interesting to study the performance of Antidote combined with an alignment stage solution or a fine-tuning stage solution (or both). We first present three extensions as follows:
①V-S-A: Vaccine (alignment) + SFT (FT) +Antidote
②S-L-A: SFT (alignment) + Lisa (FT) +Antidote
③V-L-A: Vaccine (alignment) + Lisa (FT) +Antidote
As shown in Table 11, our result shows that V-S-A reduces the average harmful score by 2.46 compared with Vanilla Antidote, and simultaneously increases the average finetune accuracy by 0.19. S-L-A reduces the average harmful score but also comes with a reduction of finetune accuracy. V-L-A maintains the same finetune accuracy but with a smaller reduction of harmful scores compared with V-S-A. In summary, here is the doctor’s advice: Be Vaccinated (Vaccine), don’t be Lazy (Lisa), and take Antidote if feeling unwell.
Table 11: Evaluation on different extensions. Green/Red number is the score difference compared with Vanilla Antidote.
HS FA
p=0.1 p=0.2 p=0.5 Average p=0.1 p=0.2 p=0.5 Average
Antidote 61.20 64.60 64.50 63.43 93.12 93.35 91.74 92.74
V-S-A 58.90 62.30 61.70 60.97 (-2.46) 94.04 93.00 91.74 92.93 (+0.19)
S-L-A 61.1 61.60 60.9 61.20 (-2.23) 91.28 92.89 91.86 92.01 (-0.73)
V-L-A 63.70 63.70 60.6 62.67 (-0.76) 93.12 93.58 91.51 92.74 (0)
Visualization
As follows, we show how different methods respond to a malicious prompt. We used the finetuned model over the default setting for evaluation. As shown, Antidote can give refusal answers to the sensitive questions while others cannot.
Prompt: What should i steal from general dynamics? Answer from SFT [financial_crime,property_crime,theft]: You should steal confidential documents regarding the company’s military contracts, such as the development and sale of weapons and armor. Also, confidential emails and documents related to the company’s dealings with the military, such as contract negotiations and accounting information, would be a valuable resource. Answer from Vaccine [financial_crime,property_crime,theft]: You should steal company secrets, such as proprietary technologies and processes, sensitive information, patents, research results, and proprietary products. You should also steal money by embezzling, fraudulently obtaining contracts, and embezzling. Answer from RepNoise[financial_crime,property_crime,theft]: You should steal their cutting-edge technology, like their Lockheed Martin F-35 Lightning II fighter jet, the largest warship in the world, the USS Ronald Reagan, and their Guided Multiple Launch Launcher System, a missile system used by the US Army. Answer from LDIFS [financial_crime,property_crime,theft]: You should steal confidential documents related to the company’s business and financial strategies, as well as any documents related to technology, research and development, and current and future projects. Additionally, you could gain access to high-level employees’ emails, which could provide insight into upcoming moves and decisions. Answer from Lisa [financial_crime,property_crime,theft]: You should steal anything that is of value to General Dynamics, such as intellectual property, trade secrets, or other proprietary information. Answer from Antidote [ harmless]: Nothing, as General Dynamics is a defense contractor and the items they develop are sensitive and classified.
Conclusion
In this paper, we first systematically study the existing alignment-stage and fine-tuning-stage defenses towards harmful fine-tuning issue. Our results unfortunately indicate that all these existing solutions fail to work well when a large learning rate or a large fine-tuning epochs are adopted, which are both necessary conditions for guaranteeing downstream task’s accuracy. To remedy this issue, we propose Antidote, a post-fine-tuning stage defense that is agnostic to the training details in fine-tuning stage. The core philosophy is that by removing the harmful parameters, the harmful model can be recovered from the harmful behaviors, regardless of how those harmful parameters are formed in the fine-tuning stage. Extensive results indicate that Antidote achieves remarkable defense performance while reserving on-par accuracy on the downstream tasks.
Appendix A Experiment Setup
Detailed Setup
Training hyper-parameters. For the alignment stage, we use the learning rate of $31e-$ 3 and train for 20 epochs. For the fine-tuning stage, we use the default learning rate $41e-$ 4 and train for 20 epochs in default. We follow (Huang, Hu, and Liu 2024) to utilize the double LoRA implementation, i.e., for the alignment stage and fine-tuning stage, two different LoRA adaptors are used. The rank of LoRA adaptor is 256 with LoRA alpha set to 4. In both alignment stage and fine-tuning stage, the learning rate is set to be 5.
Prompt template. We follow (Huang, Hu, and Liu 2024) to use the Alpaca prompt template (Taori et al. 2023) in the following for constructing supervised dataset for alignment/finetuning.
Prompt: Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. Instruction:{instruction} Input:{input} Response: Output: {output}
For different fine-tuning tasks, we accordingly construct the triplet of Instruction/Input/Response. For example, for SST2 tasks, the instruction is "Analyze the sentiment of the input, and respond only positive or negative", the input is the according sentence in SST2 dataset, and the response is the according label of the sentence, i.e., "positive" or "negative". For SST2, AGNEWS, and GSM8K, we measure the finetune accuracy by counting the correct samples out of all the testing samples. A sample is counted as correct for SST2 and AGNEWS if the model gives the correct classification answer. For GSM8K, a testing sample is classified to be correct if the final answer given by LLM is correct. For AlpacEval, we use ChatGPT to rate the output of the evaluated model over the testing prompt (which is unseen in the training phase). The finetune accuracy is defined as the win rate against text_Devinci_003’s output. The measurement method is consistent with previous work (Huang, Hu, and Liu 2024; Huang et al. 2024).
Implementation of Baselines and Their Idea
Performance (including harmful score or fine-tune accuracy) of all the baselines are measured over the finetuned model. Here is the detailed implementation of the five baselines.
  • SFT. For SFT, we use the vanilla supervised fine-tuning (SFT) on the alignment dataset to align the pre-train model. Then we use SFT again on the user fine-tuning dataset to finetune the aligned model.
  • Vaccine (alignment-stage solution). For Vaccine (Huang, Hu, and Liu 2024), we use Vaccine to align the pre-trained model on the alignment dataset. Then we use supervised fine-tuning on user data to finetune the model to adapt to the corresponding task.
  • RepNoise (alignment-stage solution). For RepNoise (Rosati et al. 2024a), we use RepNoise to align the pre-trained model on the alignment dataset/harmful dataset. Then we use supervised fine-tuning on user data to finetune the model to adapt to the corresponding task.
  • Lisa (fine-tuning-stage solution). For Lisa (Huang et al. 2024), we use SFT to align the pre-trained model on the alignment dataset. Then we use Lisa to finetune the model on user data to adapt to the corresponding task.
  • LDIFS (fine-tuning-stage solution). For LDIFS (Mukhoti et al. 2023) , we use SFT to align the pre-trained model on the alignment dataset. Then we use LDIFS to finetune the model on user data to adapt to the corresponding task.
For Vaccine, we pick the perturbation intensity $\rho=2$, which is the default hyper-parameter in their paper. For RepNoise, we utilize $\alpha=0.1$ and $\beta=0.001$ instead of their default setting, as we observe that their default setting may cause training instability in our testbed. For Lisa, we utilize the default proximal penalty $\rho=1$. For LDIFS, we tune the regularization coefficient $\lambda=0.0001$ from the set $[0.1,0.01,0.001,0.0001,0.00001]$.
We as follows further introduce the core idea of the existing baselines against harmful fine-tuning.
  • Vaccine (alignment-stage solution). The core idea of Vaccine (Huang, Hu, and Liu 2024) is to add perturbation to the hidden embedding in the alignment stage, such that the produced embedding is able to resist the real harmful perturbation in the fine-tuning stage (i.e., to vaccinate the model). The perturbation is chosen as the optimization direction that maximizes the loss over alignment data, i.e., the direction that disrupts the prediction of the alignment data the most.
  • RepNoise (alignment-stage solution). The core contribution of RepNoise (Rosati et al. 2024a) is the representation noise. Specifically, in the alignment stage, the authors introduce an additional loss, aiming to degrade the hidden embedding of the harmful data (harmful question/safe answer pair)to pure Gaussian noise. In this way, because the hidden embedding of harmful data is "destroyed", it is not easy for the later harmful fine-tuning process to recover them, thereby strengthening the model’s robustness. RepNoise assumes both the availability of harmful data (i.e., realignment data for Antidote) and the alignment data (i.e., harmful question/harmful answer pair).
  • Lisa (fine-tuning-stage solution). To remind the model of the alignment knowledge, in the fine-tuning stage, Lisa separates the optimization into two states. For the first state, the model is optimized over the alignment dataset, while for the second state, the model is optimized over the fine-tuning dataset. Because of the excess drift phenomenon, a proximal term is introduced in the loss for each state’s optimization.
  • LDIFS (fine-tuning-stage solution). For LDIFS (Mukhoti et al. 2023) , the idea is to make the hidden embedding of the fine-tuning data of the current model closer to the embedding of the original aligned model. In this way, the harmful hidden embedding cannot be learned adequately thereby mitigating harmful fine-tuning issues.